(1 point) Starting from the point (4,3,2)(4,3,2) reparametrize the curve

r(t)=(4+3t)i+(3−3t)j+(2−2t)kr(t)=(4+3t)i+(3−3t)j+(2−2t)k

in terms of arclength.
Answer #1

We reparametrization the given curve in terms of arclength with starting point (4,3,2).Non Given curve 6(+) = (4+3+) i +(3-34)j +(2-24) — (1) starting point (4,3,2) When t=o then n(o)= 41+3;+ 2K = plo) - (4,3,2)Sot 19+9+4 do Set V22 de = 522 st da = 122 []. = 122 (t-0) = t122 (+) = 4122 .: = S - $122 Now Substitute to hear (1) and get